### v3

@@ 7 Conclusion
This SLR provides an in-depth analysis of SATD detection approaches over the past decade, shedding light on the evolution of techniques from traditional keyword-based methods to sophisticated machine learning and deep learning models. Among the key findings, it is evident that while NLP-based techniques were foundational in early SATD detection, they struggle with recall and overall performance in comparison to ML and DL approaches. Transformer-based models, particularly BERT, emerged as the most effective in capturing complex patterns in textual artifacts, offering superior performance across multiple metrics.
However, several challenges persist, including issues related to dataset diversity, the scalability of models, and the explainability of AI techniques. The reliance on open-source datasets in many studies limits the generalizability of the findings, and more research is needed to adapt these models to varied industrial environments. Moreover, the imbalance of SATD datasets continues to affect the robustness of model performance, and integrating multiple textual artifacts such as commit messages and pull requests presents further complexity. Furthermore, the lack of widely adopted tools for SATD detection in industry highlights a gap between research advancements and practical applications.
In conclusion, while the field of SATD detection has advanced considerably, there remains a need for further research and collaboration between academia and industry to ensure these advancements translate into practical tools that enhance software maintainability and reduce the long-term impact of technical debt.
@@
[…]
@@
*Acknowledgement This work was financially supported by the Indonesian Education Scholarship (BPI) from the Center for Financing of Higher Education (BPPT) and Indonesia Endowment Fund for Education (LPDP).

### v4

@@ 7 Conclusion
This SLR provides an in-depth analysis of SATD detection techniques over the past decade, shedding light on the evolution of techniques from traditional keyword-based methods to sophisticated ML, DL, and Transformer-based models. Among the key findings, it is evident that while heuristic-based techniques were foundational in early SATD detection, they struggle to achieve comparable recall and overall performance to ML and DL techniques. Transformer-based models, particularly BERT, have emerged as the most effective for capturing complex patterns in textual artifacts, achieving superior performance across multiple metrics.
However, several challenges persist, including computational costs, model scalability, and the explainability of AI techniques. The reliance on open-source datasets in many studies limits the generalizability of findings, and further research is needed to adapt these models to diverse industrial environments. Moreover, the imbalance in SATD datasets continues to affect model robustness, and integrating multiple textual artifacts, such as commit messages and pull requests, introduces further complexity. Furthermore, the lack of widely adopted tools for SATD detection in the industry highlights a gap between research advancements and practical applications.
In conclusion, while the field of SATD detection has advanced considerably, further research and collaboration between academia and industry are needed to ensure that these advancements translate into practical tools that enhance software maintainability and reduce the long-term impact of technical debt.
@@
*ORCID Edi Sutoyo https://orcid.org/0000-0002-8413-5070
Andrea Capiluppi https://orcid.org/0000-0001-9469-6050
@@
*Acknowledgement This work was financially supported by the Indonesian Education Scholarship (BPI) from the Center for Higher Education Funding and Assessment (PPAPT), Indonesia Endowment Fund for Education (LPDP), and the Ministry of Higher Education, Science, and Technology of the Republic of Indonesia.